$ MC AC $56 5250 44 AVC 38 Demand MR 601 85 70 100 Sweaters per Week What is the maximum profit the firm below could earn?
Answer #1

Profit maximizing condition is when MR=MC.

Therefore, profit maximizing output is 60 sweaters and price is $50.

At this output level ,AC= $56. This implies total cost= (AC)(Q)= (56)(60)= $3360.

And total revenue = (P)(Q)= (50)(60)= $3000.

Maximum profit = (TR-TC)= $(3000-3360)= -$360 (i.e loss).